OpenStudy (anonymous):

Find dy/dx for the equation x^3-2x^2+4xy^2=20 I was instructed to use the product rule to solve, but Im confused how to actually do that. I know the product rule is f(x)g'(x)+g(x)f'(x) i'm just not really sure how to apply it to the given equation.

OpenStudy (anonymous):

sorry i type the equation worn its x^3-2x^2y+4xy^2=20

OpenStudy (anonymous):

wrong*

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}(x^3-2x^2y+4xy^2)=\frac{dy}{dx}(20)\\\frac{dy}{dx}(x^3)-2\frac{dy}{dx}(x^2y)+4\frac{dy}{dx}(xy^2)=0\)

OpenStudy (zzr0ck3r):

with me?

OpenStudy (jdoe0001):

hmmm well... 3rd term there, uses the product rule, yes the others do not

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}(x^3-2x^2y+4xy^2)=\frac{dy}{dx}(20)\\\frac{dy}{dx}(x^3)-2\frac{dy}{dx}(x^2y)+4\frac{dy}{dx}(xy^2)=0\\\frac{dy}{dx}x^3-2(\frac{dy}{dx}(x^2)y+\frac{dy}{dx}(y)x^2)+4(\frac{dy}{dx}(x)y^2+\frac{dy}{dx}(y^2)x)=0\)

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}(x^3-2x^2y+4xy^2)=\frac{dy}{dx}(20)\\\frac{dy}{dx}(x^3)-2\frac{dy}{dx}(x^2y)+4\frac{dy}{dx}(xy^2)=0\\\frac{dy}{dx}x^3-2[\frac{dy}{dx}(x^2)y+\frac{dy}{dx}(y)x^2]+4[\frac{dy}{dx}(x)y^2+\frac{dy}{dx}(y^2)x]=0\)

OpenStudy (zzr0ck3r):

i use the product rule inbetween the []

OpenStudy (zzr0ck3r):

am I talking to my self?

OpenStudy (anonymous):

No i'm just trying to go through it and see if I'm picking up on it

OpenStudy (zzr0ck3r):

word

OpenStudy (anonymous):

I know this is a really stupid question but do you have to put the dy/dx in front of each number like you did ?

OpenStudy (zzr0ck3r):

not really, you could just take the derivative, but I figured this would help you see exactly what we are doing

OpenStudy (anonymous):

okay thanks

OpenStudy (zzr0ck3r):

now since we are doing implicit differentiation we have \(\frac{dy}{dx}(x^3-2x^2y+4xy^2)=\frac{dy}{dx}(20)\\\frac{dy}{dx}(x^3)-2\frac{dy}{dx}(x^2y)+4\frac{dy}{dx}(xy^2)=0\\\frac{dy}{dx}x^3-2[\frac{dy}{dx}(x^2)y+\frac{dy}{dx}(y)x^2]+4[\frac{dy}{dx}(x)y^2+\frac{dy}{dx}(y^2)x]=0\\ 3x^2-2[2xy+\frac{dy}{dx}x^2]+4[y^2+2y\frac{dy}{dx}x]=0\\3x^2-4xy-2\frac{dy}{dx}x^2+4y^2+8yx\frac{dy}{dx}=0\)

OpenStudy (zzr0ck3r):

now solve for \(\frac{dy}{dx}\)